A bond with a face value of $1,000 has 10 years until maturity, carries a coupon rate of 7.5%, and sells for $1,150. Interest is paid annually. (Assume a face value of $1,000 and annual coupon payments.)

a. If the bond has a yield to maturity of 10.5% 1 year from now, what will its price be at that time? (Do not round intermediate calculations. Round your answer to nearest whole number.)

b. What will be the rate of return on the bond? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. Negative amount should be indicated by a minus sign.)

c. If the inflation rate during the year is 3%, what is the real rate of return on the bond? (Assume annual interest payments.) (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. Negative amount should be indicated by a minus sign.)
Answer #1

Price=coupon rate*par value/yield to maturity*(1-1/(1+yield to maturity)^years to maturity)+par value/(1+yield to maturity)^years to maturity=7.5%*1000/10.5%*(1-1/1.105^9)+1000/1.105^9=830.6102837

Rate of return=(Price 1 year later+coupon rate*par value)/Price now-1=(830.6102837+7.5%*1000)/1150-1=-21.251%

Real return=(1+nominal return)/(1+inflation)-1=(1-21.251%)/(1+3%)-1=-23.545%